# Rad4XCNN: a new agnostic method for post-hoc global explanation of CNN-derived features by means of radiomics

### v1

@@ 3 Materials and Methods
Figure 1 shows the flow diagram of the proposed method. For the deep workflow ResNet50, DenseNet121, and ViT-B32 were pre-trained on the BUSI dataset. Then the extracted deep features were employed both for classification (via multilayer perception) and for explanation (local explanation via visualization maps). For the radiomic workflow, radiomic features were extracted, preprocessed, and used to train shallow classifiers. Radiomic and deep features were used as input for Rad4XCNN providing a global explanation.
@@ 3.1 Datasets Description
Breast Ultrasound Images Dataset (BUSI)
The Breast Ultrasound Images Dataset (BUSI) Al-Dhabyani et al. (2020) comprises 780 US images obtained from 600 patients aged 25-75 years in 2018, alongside corresponding masks in PNG format. The images - containing the Focal Breast Lesions (FBLs) - measure around $500\times 500$ pixels. They are categorized into three classes: benign, malignant, and normal. Among these, 437 images present benign lesions, 210 are malignant, and 133 are classified as normal. Normal samples were not used in the study. The dataset exhibits class imbalance, with 67.49% benign samples and 32.51% malignant.
Palermo and Cefalù Datasets
Two in-house datasets were collected on two centers, referred to as Palermo and Cefalù datasets. Considering the Palermo site, 237 breast cancer patients were enrolled in 2021 at Policlinico Universitario “P. Giaccone”. The obtained dataset was split into two subsets (with an 80/20 ratio) used for the training and testing phases, respectively. Instead, considering the Cefalù site, 115 breast cancer patients were enrolled in 2022 at Fondazione Istituto "G. Giglio" Breast Unit. This dataset was used for model external validation. Table 1 provides a detailed explanation of the in-house datasets.
FBL images of both datasets were acquired through the B-mode US modality. In particular, two experienced breast radiologists (one for each center, more than 30 years of experience in breast imaging) performed US acquisition through two identical ultrasound RS85 (manufactured by Samsung Medison, Co. Ltd.) units, each provided with the same 3–12 MHz linear transducer. Acquired FBLs were labeled according to US BI-RADS criteria Sickles and D’Orsi (2014) by expert radiologists and successively segmented.
In both centers, indications for breast US included:
  • a palpable mass detected on physical examination;
  • dense breasts;
  • detected lesions from adjunct mammography examination;
  • patients with mastodynia;
  • young patients having a family history
  • patients in a follow-up for benign breast nodules
Patients have not undergone any intervention or surgery on lesions before US examination. Exclusion criteria included lack of adequate standard of reference (refusal to undergo biopsy or inconsistent follow-up).
@@ 3.2 Deep Architecture Training
This section includes the workflow description to train deep architectures. In particular, preprocessing strategies, data preparation for training and test, and the implemented architectures are discussed. In addition, state-of-the-art methods for deep features explanation were exposed as well as our novel Rad4XCNN method.
Image preprocessing
Patches were extracted from the whole image using the available masks. All patches were resized to size $128\times 128$, specifically: i) images with a size larger than $128\times 128$ were subsampled, and ii) zero-padding was applied to images with a smaller size. Images were normalized before model training.
Training and test protocol
The BUSI dataset was divided into training and test sets, using the ratio 80:20. The training set was split again into training and validation sets, with the same ratio. In addition, the training set was balanced by considering random flip, rotation, contrast enhancement, translation, and zoom. Considering the recent criticism about the BUSI dataset Pawłowska et al. (2023), it was used only to obtain an optimized pre-trained model. Regarding the Palermo in-house dataset, 15% was used as an independent test set, while the remaining 85% was used to implement a 5-fold cross-validation (CV). The Cefalù dataset was used only for external testing. The best model in terms of accuracy within the CV procedure was selected for the Palermo internal test and the Cefalù external test.
Regarding the architectures, state-of-the-art models were employed, including Densenet121, Resnet50, and ViT-B32. The extracted feature vector in ResNet50 and DenseNet121 is 2048 and 1024 neurons, respectively. Their head was implemented using (1) a dropout layer, (2) a dense layer with 512 neurons, (3) another dropout, and (4) the last dense layer with 2 neurons. Regarding ViT-B32, the extracted feature vector was 768 dimensions, followed by (1) a dropout layer, (2) a batch normalization layer, and (3) two dense layers of 512 and 2 neurons, respectively. The softmax activation function was used on the last dense layer and the ReLU for the others.
Saliency Maps Computation
@@
Here $\frac{1}{z}\sum_{i}\sum_{j}$ represents the global average pooling operator over the width and height dimensions (indexed by $i$ and $j$ respectively) and $\frac{\partial y^{c}}{\partial A^{k}_{ij}}$ gradients via backpropagation. The weights $\alpha^{c}_{k}$ represent a partial linearization capture the importance of feature map k for a target class c. Grad-CAM’s dependence on class $c$ makes it a class discriminative method. Although Grad-CAM appears as one of the most widely used approaches, has two main limitations related to the gradient computation: i) saturation: Gradient can vanish due to the saturation problem for Sigmoid function or the zero-gradient region of ReLU function; and ii) false confidence: Activation maps with higher weights show lower contributions to the network’s output compared to a zero baseline Wang et al. (2020). In addition, backpropagating any quantity requires additional computational overhead and assumes that classifiers produced correct decisions, and whenever a wrong decision is made, all mentioned methods will produce wrong or distorted visualizations Muhammad and Yeasin (2020).
In contrast to Grad-CAM Selvaraju et al. (2017), which use the gradient information flowing into the last convolutional layer to represent the importance of each activation map, the Channel-wise Increase of Confidence (CIC) was introduced in Score-CAM Wang et al. (2020).
@@
where $C(\cdot)$ denotes the Channel-wise Increase of Confidence (CIC) defined to measure the importance of each activation map.
EigenCAM is a gradient-free method and uses the principal components from the extracted feature maps Muhammad and Yeasin (2020). For this reason, it can overcome the problem of distorted displays in case of incorrect predictions. Let $I$ represent the input image of size $(i\times j)$, $I\in\mathbb{R}^{i\times j}$, and let $W_{L=k}$ represent the combined weight matrix of the first $k$ layers of size $(m,n)$. The class activated output is the image $I$ projected onto the last convolution layer $L=k$ and is given by $O_{L=k}=W^{T}_{L=k}I$, factorizing $O_{L=k}$ using singular value decomposition to compute the principal components of $O_{L=k}$ gives $O_{L=K}=U\Sigma V^{T}$. $U$ is a $M\times M$ orthogonal matrix and the column $U$ are the left singular vectors, $\Sigma$ is a diagonal matrix of size $M\times N$ with singular values along the diagonal and $V$ are the left singular vectors. The class activation map, $L_{EigenCAM}$ is given by the projection of $O_{L=k}$ on the first eigenvector:
@@
where $V_{1}$ is the first eigenvector in the V matrix. Eigen-CAM is class-independent, e.g. the result of saliency maps is not tied to any class.
@@ 3.3 Radiomic Models Training
Radiomic workflow includes ROI segmentation, feature extraction, feature preprocessing, selection, and model training.
ROI Segmentation
ROI segmentation was performed by radiologists from each B-mode image of the breast lesion through S-Detect™. This software is commercially available, licensed for clinical use, and installed on the RS85 ultrasound unit.
Feature Extraction
A total of 474 radiomic features were extracted using PyRadiomics Van Griethuysen et al. (2017), a toolkit compliant with the IBSI Zwanenburg et al. (2020). IBSI aims to standardize the radiomic analysis process, improving workflow reproducibility. Extracted features belong to the following categories: shape 2D; first-order intensity histogram statistics; Gray Level Co-occurrence Matrix features (GLCM); Gray Level Run Length Matrix (GLRLM); Gray Level Size Zone Matrix (GLSZM); Gray Level Dependence Matrix (GLDM); Neighboring Gray Tone Difference Matrix (NGTDM). Features were extracted discretizing images to 255 gray levels. In addition, considering the predictive capability of wavelet-derived radiomic features compared to the original ones Prinzi et al. (2023b), non-shape features were extracted also considering the Haar wavelet transform.
Features Preprocessing and Selection
Non-redundant and informative features were selected Militello et al. (2022); Papanikolaou et al. (2020) through: i) near-zero variance analysis to remove features with low information content (variance cutoff $\sigma^{2}=0.005$); ii) correlation analysis to remove highly correlated features (the Spearman correlation with a 0.9 threshold value was considered). The Sequential Feature Selector (SFS) algorithm combined with XGBoost (XGB), Random Forest (RF), and SVM classifiers was used to select the most discriminating features. Accuracy was used as a metric to maximize.
Predictive Model Setup
This study employed three distinct shallow classifiers: RF, XGB, and SVM. These algorithms have proven suitable for tabular data in small dataset scenarios Prinzi et al. (2022). RF and XGB were trained using 100 estimators. For XGB a learning rate of 0.3 and gain as the importance type were employed. SVM training involved standardizing the features and RBF kernel. The predictive model setup was performed on the training set using the XGB classifier and considering a stratified 10-fold Cross-Validation procedure repeated 20 times. The best model in terms of accuracy found during the 20-repeated 10-fold cross-validation procedure was selected for the test.
@@ 3.4 Deep Features Explanation by means of Radiomics
Rad4XCNN aims to explain the deep features extracted through the intrinsic meaning of radiometric features. Let $f_{r}=(r_{1},r_{2},...,r_{a})$ denote the radiomic features, and $f_{d}=(d_{1},d_{2},...,d_{b})$ denote the deep features, where $a$ and $b$ are the number of radiomic and deep features, respectively. The method assumes the use of a correlation index, and Sperman’s test was employed in this work. The Spearman correlation coefficient between $f_{r}$ and $f_{d}$ is given by:
@@ 4 Experimental Results
Deep models were trained using Adam optimizer and 8 as batch size. For ResNet50 and DenseNet121 $10^{-4}$ was used as the learning rate, and $10^{-5}$ for ViT. A rate of 0.5 was used for dropout layers. The BUSI dataset was employed to obtain an initial pre-trained model, and 80% of the Palermo dataset was used for fine-tuning. Then, the remaining 20% of the Palermo dataset was used as an internal test and the whole Cefalù dataset for an external test.
@@ 4.1 Deep and Radiomic Models Performance
Internal and external test performance are shown in Table 3, in which again the superiority of deep models over radiomic-based approaches is shown. In the Palermo dataset, ResNet achieves an accuracy of $72.22\%$ and AUROC of $0.733$, while DenseNet and ViT achieve even higher accuracies and AUROC scores. These results highlight the effectiveness of deep learning methodologies in extracting meaningful features from the datasets, leading to improved diagnostic performance. On the Cefalù dataset, a similar trend is observed, with deep models outperforming radiomic-based approaches. In particular, ResNet proved the best generalization capability on the external test, showing an accuracy of $73.04\%$ and a good balance between sensitivity and specificity. It is important to emphasize that although the performance of ResNet on the internal test is much lower than ViT, on the external test ResNet maintains its performance, as opposed to ViT which achieves a heavy degradation. For this reason, we consider ResNet the best trained model, and further considerations on explainability are provided for ResNet.
In contrast, models incorporating radiomic features combined with traditional machine learning algorithms consistently demonstrate inferior performance compared to deep learning models on both internal and external test datasets. These results underscore the limitations of radiomic-based approaches in capturing complex patterns and features present in medical images, particularly when compared to the feature extraction capabilities of deep learning architectures.
{adjustwidth}
-0.1cm Model Accuracy AUROC Sens Spec PPV NPV ResNet50 $93.62\pm 9.19$ $0.95\pm 0.09$ $91.20\pm 14.83$ $96.26\pm 4.6$ $95.23\pm 5.91$ $93.01\pm 12.00$ DenseNet121 $93.17\pm 13.66$ $0.94\pm 0.12$ $93.33\pm 13.33$ $93.00\pm 14.00$ $93.33\pm 13.33$ $93.00\pm 14.00$ ViT-B32 $93.65\pm 11.5$ $0.97\pm 0.6$ $92.16\pm 12.95$ $95.00\pm 10.00$ $94.74\pm 10.53$ $92.80\pm 12.42$ Rad + XGB $61.49\pm 9.88$ $0.65\pm 0.10$ $53.63\pm 16.79$ $67.53\pm 14.13$ $56.63\pm 13.19$ $66.05\pm 9.59$ Rad + RF $59.93\pm 9.19$ $0.60\pm 0.11$ $39.61\pm 14.31$ $75.35\pm 13.01$ $56.66\pm 17.76$ $62.10\pm 7.34$ Rad + SVM $65.05\pm 8.31$ $0.62\pm 0.10$ $55.83\pm 15.05$ $72.09\pm 12.13$ $61.52\pm 10.06$ $68.70\pm 8.60$
{adjustwidth}
-0.7cm Palermo Dataset Cefalù Dataset Model Acc AUROC Sens Spec PPV NPV Acc AUROC Sens Spec PPV NPV ResNet50 $72.22$ $0.733$ $66.67$ $76.19$ $66.67$ $76.19$ $73.04$ $0.757$ $77.78$ $70.00$ $62.5$ $83.05$ DenseNet121 $77.78$ $0.784$ $73.33$ $80.95$ $73.33$ $80.95$ $70.44$ $0.755$ $68.89$ $71.43$ $60.78$ $78.13$ ViT-B32 $83.33$ $0.848$ $93.33$ $76.19$ $73.68$ $94.12$ $62.61$ $0.742$ $77.78$ $52.86$ $51.47$ $78.72$ Rad + XGB $47.22$ $0.43$ $46.15$ $47.82$ $33.33$ $47.22$ $61.11$ $00.60$ $60.86$ $61.35$ $62.89$ $60.86$ Rad + RF $58.33$ $0.61$ $61.15$ $56.52$ $44.44$ $58.33$ $51.12$ $0.55$ $54.78$ $53.47$ $56.25$ $54.78$ Rad + SVM $58.33$ $0.58$ $53.84$ $60.86$ $43.75$ $58.33$ $58.17$ $0.62$ $58.26$ $58.08$ $60.09$ $58.26$
Table 3: Internal and external test performance using ResNet, DenseNet, ViT, and radiomic models on the Palermo and Cefalù test datasets.

### v2

@@ 3 Materials and Methods
Figure 1 shows the flow diagram of the proposed method. For the deep workflow ResNet50, DenseNet121, and ViT-B32 were pre-trained on the BUSI dataset. Then transfer learning was implemented using the two in-house datasets. The extracted deep features were employed both for classification (via multilayer perceptron) and for explanation (local explanation via visualization maps). For the radiomic workflow, radiomic features were extracted, preprocessed, and used to train shallow classifiers. Radiomic and deep features were used as input for Rad4XCNN providing a global explanation.
@@ 3.1 Datasets Description
Breast Ultrasound Images Dataset (BUSI) -
The BUSI Al-Dhabyani et al. (2020) comprises 780 ultrasound images obtained from 600 patients aged 25-75 years in 2018, alongside corresponding masks in PNG format. The images, containing the Focal Breast Lesions (FBLs) and measuring $500\times 500$ pixels, are categorized into three classes: benign, malignant, and normal. Among these, 438 images present benign lesions, 211 are malignant, and 133 are classified as normal. Normal samples were not used in the study. The dataset exhibits class imbalance, with 67.49% (438 images) benign and 32.51% (211 images) malignant samples.
Table 1 provides the details concerning the number of images available of the BUSI dataset before and after the data augmentation, which was performed to increase the training set of the minority class: only the malignant class was augmented. Moreover, the table reports the dataset subdivision for training, validation, and test phases. After the initial 80:20 ratio between ‘Training/Validation’ and ‘Test’, the ‘Training/Validation’ set was again divided by applying an 80:20 ratio to obtain the ‘Training’ and ‘Validation’ sets.
Table 1: Details concerning the number of images available of the BUSI dataset before and after the data augmentation, which was performed to increase the training set of the minority class. Moreover, the table reports the dataset subdivision for training, validation, and test phases. (*) indicates the number of images obtained after data augmentation.
Lesion Class Training/Validation (80%) Test (20%)
Training (80%) Validation (20%)
before data augmentation after data augmentation
Benign 280 280 70 88
Malignant 135 270 (∗) 34 42
Total 415 550 104 130
@@
Table 3: Details on proprietary datasets: splits employed for model training and test.
Lesion Class Palermo Dataset Cefalù Dataset
Total Training Test Test
Benign 132 106 26 70
Malignant 105 84 21 45
Total 237 190 47 115
Palermo and Cefalù Datasets -
Two in-house datasets were collected on two centers, referred to as Palermo and Cefalù datasets. Considering the Palermo site, 237 breast cancer patients were enrolled in 2021 at Policlinico Universitario “P. Giaccone”. This dataset was employed for model tuning and internal test. Instead, considering the Cefalù site, 115 breast cancer patients were enrolled in 2022 at Fondazione Istituto "G. Giglio" Breast Unit. This dataset was used only for model external validation. Table 2 provides a detailed explanation of the in-house datasets. FBL images of both datasets were acquired through the B-mode ultrasound modality. Two expert breast radiologists (one from each institution with over 30 years of expertise in breast imaging) used two identical ultrasound RS85 (Samsung Medison, Co. Ltd.) devices, each with a 3–12 MHz linear transducer. Acquired FBLs were labeled using the ultrasound BI-RADS criteria Sickles and D’Orsi (2014), and successively the images were segmented. The adopted inclusion criteria for our breast ultrasound in-house datasets are discussed here Bartolotta et al. (2021). Patients had not received any intervention or surgery on lesions before the ultrasound examination. Lack of biopsy or irregular follow-up were the exclusion criteria Bartolotta et al. (2024). Table 3 shows the split of the proprietary datasets used for training and testing. A cross-validation was implemented using the training set.
@@ 3.2 Deep Architecture Training
This section includes a description of the workflow used to train deep architectures. In particular, preprocessing strategies, data preparation for training and test, and the implemented architectures are discussed. In addition, state-of-the-art methods for deep features explanation were exposed as well as our novel Rad4XCNN method. The diagram reported in Figure 2 depicts the details about the training, the testing, and the external validation related to deep workflow.
Image preprocessing -
Tumor patches were extracted from the whole image using the available masks. All patches were resized to size $128\times 128$, specifically: i) images with a size larger than $128\times 128$ were subsampled, and ii) zero-padding was applied to images with a smaller size. Images were normalized before model training.
Training and test protocol -
Using an 80:20 ratio, the BUSI dataset was split into training and test sets. With the same ratio, the training set was divided once more into training and validation sets. Whereas having unbalanced datasets can lead to models affected by bias Hasib et al. (2020), the training set was balanced by considering random flip, rotation, contrast enhancement, translation, and zoom. The purpose of the BUSI dataset, in light of the recent criticism Pawłowska et al. (2023), was to produce an optimized pre-trained model solely. Regarding the Palermo dataset, 20% was used as an independent test set, while the remaining 80% was used to implement a 5-fold cross-validation (CV). The Cefalù dataset was used only for external validation. The best model in terms of accuracy within the CV procedure was selected for the Palermo internal test and the Cefalù external validation.
Saliency Maps Computation -
@@
Here $\frac{1}{z}\sum_{i}\sum_{j}$ represents the global average pooling operator over the width and height dimensions (indexed by $i$ and $j$ respectively) and $\frac{\partial y^{c}}{\partial A^{k}_{ij}}$ gradients via backpropagation. The weights $\alpha^{c}_{k}$ reflect a partial linearization and quantify feature map $k$ importance for a given target class $c$. Grad-CAM’s dependence on class $c$ makes it a class discriminative method. Although Grad-CAM appears as one of the most widely used approaches, has two main limitations related to the gradient computation: i) saturation: The zero-gradient area of the ReLU function or the saturation problem for the Sigmoid function may cause the gradient to vanish; and ii) false confidence: when compared to a zero baseline, activation maps with larger weights exhibit smaller contributions to the network’s output Wang et al. (2020). Furthermore, backpropagating any quantity has extra computing overhead and relies on the assumption that classifiers generated accurate prediction; in the event that an incorrect decision is made, all of the aforementioned techniques will result in inaccurate or distorted representations Muhammad and Yeasin (2020).
Score-CAM introduced the Channel-wise Increase of Confidence (CIC), in contrast to Grad-CAM Selvaraju et al. (2017), which uses the gradient information coming into the last convolutional layer to reflect the relevance of each activation map Wang et al. (2020).
@@
where the CIC, used to gauge each activation map’s significance, is shown by the symbol $C(\cdot)$.
EigenCAM is a gradient-free method and uses the principal components from the extracted feature maps Muhammad and Yeasin (2020). For this reason, it can overcome the problem of distorted displays in case of incorrect predictions. Let $W_{L=k}$ represent the combined weight matrix of the first $k$ layers of size $(m,n)$, and let $I$ represent the input image of size $(i\times j)$, $I\in\mathbb{R}^{i\times j}$. The image $I$ projected onto the last convolution layer $L=k$ is the class-activated output, and it is provided by $O_{L=k}=W^{T}_{L=k}I$. The principal components of $O_{L=k}$ can be computed by factorizing $O_{L=k}$ using singular value decomposition, which yields $O_{L=K}=U\Sigma V^{T}$. $\Sigma$ is a diagonal matrix of size $M\times N$ with singular values along the diagonal, and $V$ are the left singular vectors. $U$ is an orthogonal matrix of size $M\times M$, and the left singular vectors are found in the column $U$. The projection of $O_{L=k}$ on the first eigenvector yields the class activation map:
@@
where the first eigenvector in the V matrix is $V_{1}$. Eigen-CAM is class-independent, e.g. the result of saliency maps is not tied to any class.
Figure 2: Diagram depicting the details about the training, the testing, and the external validation related to deep and radiomic workflow.
The three methods (e.g. Grad-CAM, Score-CAM, and Eigen-CAM) were implemented to compare the different explanations provided and, thus, highlight existing gaps and the need for alternative explanation techniques.
@@ 3.3 Radiomic Models Training
Radiomic workflow includes ROI segmentation, feature extraction, feature preprocessing, selection, and model training. The diagram reported in Figure 2, already introduced in the previous section, depicts the details about the training, the testing, and the external validation related to radiomic workflow.
ROI Segmentation -
Radiologists used S-DetectTM to perform ROI segmentation from each B-mode image of the breast lesion. Installed on the RS85 ultrasound machine, this software is licensed for clinical usage and is available for purchase.
Feature Extraction -
PyRadiomics Van Griethuysen et al. (2017) is a toolkit that complies with the IBSI Zwanenburg et al. (2020), and was used to extract a total of 474 radiomic features. By standardizing the radiomic analysis procedure, IBSI increases workflow reproducibility. Extracted features belong to the following categories: shape 2D; first-order intensity histogram statistics; Gray Level Co-occurrence Matrix features (GLCM); Gray Level Run Length Matrix (GLRLM); Gray Level Size Zone Matrix (GLSZM); Gray Level Dependence Matrix (GLDM); Neighboring Gray Tone Difference Matrix (NGTDM). Features were extracted discretizing images to 255 gray levels. Furthermore, considering the high productivity of wavelet-derived radiomic features in comparison to the original ones Prinzi et al. (2023b), non-shape features were extracted also considering the Haar wavelet transform.
Features Preprocessing and Selection -
Non-redundant and informative features were selected Militello et al. (2022); Papanikolaou et al. (2020) through: i) near-zero variance analysis to eliminate features with poor information content (variance cutoff $\sigma^{2}=0.005$); ii) correlation analysis to exclude strongly correlated features (the Spearman correlation with a 0.9 threshold value was evaluated). Eventually, the Sequential Feature Selector (SFS) algorithm was employed in conjunction with the Random Forest (RF), XGBoost (XGB), and Support Vector Machine (SVM) classifiers. Accuracy was used as a metric to maximize.
Predictive Model Setup -
This study employed three distinct shallow classifiers: RF, XGB, and SVM. These algorithms have proven suitable for tabular data in small dataset scenarios Prinzi et al. (2022). RF and XGB were trained using 100 estimators. For XGB a learning rate of 0.3 and gain as the importance type were employed. SVM was trained using RBF kernel and features were standardized. A stratified 10-fold Cross-Validation approach that was repeated 20 times was taken into consideration. To perform the test phase, the model with the highest accuracy after 20 repetitions of a 10-fold cross-validation process was selected.
@@ 3.4 Deep Features Explanation by means of Radiomics
Rad4XCNN aims to explain the deep features extracted through the intrinsic meaning of radiomic features. Let $f_{r}=(r_{1},r_{2},...,r_{a})$ denote the radiomic features, and $f_{d}=(d_{1},d_{2},...,d_{b})$ denote the deep features, where $a$ and $b$ are the number of radiomic and deep features, respectively. The method assumes the use of a correlation index, and Sperman’s test was employed in this work. The Spearman correlation coefficient between $f_{r}$ and $f_{d}$ is given by:
@@
Model Acc AUROC Sens Spec PPV NPV
ResNet50 $93.62\pm 9.19$ $0.95\pm 0.09$ $91.20\pm 14.83$ $96.26\pm 4.6$ $95.23\pm 5.91$ $93.01\pm 12.00$
DenseNet121 $93.17\pm 13.66$ $0.94\pm 0.12$ $93.33\pm 13.33$ $93.00\pm 14.00$ $93.33\pm 13.33$ $93.00\pm 14.00$
ViT-B32 $93.65\pm 11.5$ $0.97\pm 0.6$ $92.16\pm 12.95$ $95.00\pm 10.00$ $94.74\pm 10.53$ $92.80\pm 12.42$
Rad + XGB $61.49\pm 9.88$ $0.65\pm 0.10$ $53.63\pm 16.79$ $67.53\pm 14.13$ $56.63\pm 13.19$ $66.05\pm 9.59$
Rad + RF $59.93\pm 9.19$ $0.60\pm 0.11$ $39.61\pm 14.31$ $75.35\pm 13.01$ $56.66\pm 17.76$ $62.10\pm 7.34$
Rad + SVM $65.05\pm 8.31$ $0.62\pm 0.10$ $55.83\pm 15.05$ $72.09\pm 12.13$ $61.52\pm 10.06$ $68.70\pm 8.60$
Table 5: Internal test and external validation performance using ResNet, DenseNet, ViT, and radiomic models on the Palermo and Cefalù test datasets.
Palermo Hospital Dataset Cefalù Hospital Dataset
Model Acc AUROC Sens Spec PPV NPV Acc AUROC Sens Spec PPV NPV
ResNet50 $72.22$ $0.733$ $66.67$ $76.19$ $66.67$ $76.19$ $73.04$ $0.757$ $77.78$ $70.00$ $62.5$ $83.05$
DenseNet121 $77.78$ $0.784$ $73.33$ $80.95$ $73.33$ $80.95$ $70.44$ $0.755$ $68.89$ $71.43$ $60.78$ $78.13$
ViT-B32 $83.33$ $0.848$ $93.33$ $76.19$ $73.68$ $94.12$ $62.61$ $0.742$ $77.78$ $52.86$ $51.47$ $78.72$
Rad + XGB $47.22$ $0.43$ $46.15$ $47.82$ $33.33$ $47.22$ $61.11$ $00.60$ $60.86$ $61.35$ $62.89$ $60.86$
Rad + RF $58.33$ $0.61$ $61.15$ $56.52$ $44.44$ $58.33$ $51.12$ $0.55$ $54.78$ $53.47$ $56.25$ $54.78$
Rad + SVM $58.33$ $0.58$ $53.84$ $60.86$ $43.75$ $58.33$ $58.17$ $0.62$ $58.26$ $58.08$ $60.09$ $58.26$
@@ 4 Experimental Results
Deep models were trained using Adam optimizer and 8 as batch size. For ResNet50 and DenseNet121 $10^{-4}$ was used as the learning rate, and $10^{-5}$ for ViT. A rate of 0.5 was used for dropout layers. The BUSI dataset was employed to obtain an initial pre-trained model, and 80% of the Palermo dataset was used for fine-tuning. Then, the remaining 20% of the Palermo dataset was used as an internal test, and the whole Cefalù dataset for external validation.
@@ 4.1 Deep and Radiomic Models Performance
Internal test and external validation performance are shown in Table 5, in which again the superiority of deep models over radiomic-based approaches is shown. In the Palermo dataset, ResNet achieves an accuracy of $72.22\%$ and AUROC of $0.733$, while DenseNet and ViT achieve even higher accuracies and AUROC scores. These results highlight the effectiveness of deep learning methodologies in extracting meaningful features from the datasets, leading to improved diagnostic performance. On the Cefalù dataset, a similar trend is observed, with deep models outperforming radiomic-based approaches. In particular, ResNet proved the best generalization capability on the external validation, showing an accuracy of $73.04\%$ and a good balance between sensitivity and specificity. It is important to emphasize that although the performance of ResNet on the internal test is much lower than ViT, on the external validation ResNet maintains its performance, as opposed to ViT which achieves a heavy degradation. For this reason, we consider ResNet the best-trained model, and further considerations on explainability are provided for ResNet.
In contrast, models incorporating radiomic features combined with traditional machine learning algorithms consistently demonstrate inferior performance compared to deep learning models on both internal test and external validation datasets. These results underscore the limitations of radiomic-based approaches in capturing complex patterns and features present in medical images, particularly when compared to the feature extraction capabilities of deep learning architectures.